RPC weekly meeting

Europe/Zurich
CERN

CERN

Davide Boscherini (Universita e INFN, Bologna (IT)), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Mohamed Zaazoua (University of Science and Technology of China (CN))
Description

Weekly meeting about the RPC system activities

Zoom Meeting ID
61128581037
Host
Davide Boscherini
Useful links
Join via phone
Zoom URL
    • 09:15 09:25
      Communication to/from gas group 10m
      Speakers: Beatrice Mandelli (CERN), Enrico Pastori (INFN e Universita Roma Tor Vergata (IT)), Gianluca Rigoletti (CERN), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Kacper Kapusniak (CERN), Roberto Guida (CERN)
    • 09:25 09:35
      Detector and trigger reports 10m
      Speakers: Claudio Luci (Sapienza Universita e INFN, Roma I (IT)), Daniela Katherinne Paredes Hernandez (University of Hong Kong (HK)), Enrico Pastori (INFN e Universita Roma Tor Vergata (IT)), Enrico Tusi (INFN e Universita Roma Tor Vergata (IT)), Eric Ballabene (University and INFN, Bologna (IT)), Dr Federico Lasagni Manghi (Universita e INFN, Bologna (IT)), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Mr Mohamed Zaazoua (University of Science and Technology of China (CN)), Sinem Simsek (Istinye University (TR))

      Eric's Shift:

      Some LV Vpd got switched off automatically due to board temp monitoring and were switched back on https://atlasop.cern.ch/elisa/display/604977

      LV with decrease Iee https://atlasop.cern.ch/elisa/display/605297

      My shift:

      DCS alarm: RPC CAEN OPC HV/LV Quadrant 1

      DPE: ATLRPCSCS:OpcServer_ATLRPCH1.watchDog 2025.07.19 09:56:08.038: FATAL

       

      Subject: ADC isuue on the RPC

      The following board went status 10 standing for 17VPwS or  ADC counter fail. This lead to a fatal in Oct1. Now the board is masked and we will try to recover it when the there is no beam. 

      Location: Y_2904_X7_C2/H1/br02/crate0/slot08, Model: A3801, Status: 10, Temp: 0, Fsm: b040070, Mask: 1, Hw: H1:br02.ch013

      About 1min before that I noticed BOS5.A02.CO.DP1.Ly0.DZ2 gap channel showing a sharp jump in the current up to 100 uA (masked and  kept off  the HV channel until detailed debug is conducted). 

      Subject: RPC recovery after lightning

      Lightning did trip 2 48V generators, then to recover needed crate crate0 HW reset which fixed all boards including the ADC which was stuck reported in the previous elog. All  HV, Vee, Vpd and Vpad channels were turned On. The remaining warnings will disappear after starting a new run. 

      Subject: resolving the fatal in the Alarm screen from yesterday "RPC caen OPC HV/LV Quadrant 4"

      Because of the fatal in the alarm screen "RPC caen OPC HV/LV Quadrant 4" (we didn't want to act during ongoing run yesterday ), a reset of the h2 is conducted and the fatal is resolved. Trigger expert checked the trigger infrastructure and  confirmed everything looks fine.

    • 09:35 09:45
      Vth tuning in noisy trigger towers 10m
      Speakers: Claudio Luci (Sapienza Universita e INFN, Roma I (IT)), Eric Ballabene (University and INFN, Bologna (IT)), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Sinem Simsek (Istinye University (TR))
    • 09:45 09:55
      Resin update 10m
      Speaker: Ahmet Renklioglu (Istinye University (TR))
    • 09:55 10:05
      Efficiency studies 10m
      Speakers: Claudio Luci (Sapienza Universita e INFN, Roma I (IT)), Daniele Truncali (INFN e Universita Roma Tor Vergata (IT)), Marco Sessa (University Federico II and INFN, Naples (IT))